ACC 302

1)      Smith Incorporated acquired a piece of equipment at a total cost of $4,200,000. They use the straight- line method for financial reporting and MACRS depreciation for tax purposes. The asset has a six-year life for book purposes and because they use half year convention the cost is depreciated over six years for tax purposes also. The tax rate is 40%. The following information is available:

 

Year

Income before   tax and Depreciation

Tax

Depreciation

GAAP

Depreciation

1

$920,000

$840,000

$700,000

2

1,600,000

1,344,000

700,000

3

1,780,000

806,400

700,000

4

2,100,000

483,840

700,000

5

1,750,000

483,840

700,000

6

1,200,500

241,920

700,000

Total


$4,200,000

$4,200,000

 

Required (14 Points)

 

a.       Determine the balance of the deferred tax account at the end of each year

b.       Prepare the journal entries to record the tax provision for each year.
Answer #1

a)Defered tax account balance

year 1=840000-700000=140,000 tax=40%*140000=56000

year 2=1344000-700000=644,000 tax=40%*644000=257600

Year 3=806400-700000=106,400 tax=40%*106400=42560

Year 4=483840-700000=-216,160 tax=40%*-216160=-86464

year 5=483840-700000=-216,160 tax=40%*-216160=-86464

Year 6=241920-700000=-458,080 tax=40%*-458080=-183232

b)profit & loss (db) 56000

defered tax laibility (cR) 56000

profit & loss (db) 257600

defered tax laibility (cR) 257600

profit & loss (db) 42560

defered tax laibility (cR) 42560

profit & loss (cr) 86464

defered tax asset (db) 86464

profit & loss (cr) 86464

defered tax asset (db) 86464

profit & loss (cr) 183232

defered tax asset (db) 183232


answered by: ANURANJAN SARSAM